OpenStudy (anonymous):

A pizza shop makes large pizzas with a target diameter of 16 inches. A pizza is acceptable if its diameter of a pizza is within 3 (times) 2 (to the power of) -2 in of the target diameter. Let d represent the diameter of a pizza. Write an inequality for the range of acceptable large pizza diameters in inches. (Where do I even start?)

OpenStudy (anonymous):

\[ |d_{target}-d_{pizza}|\leq 3\cdot2^{-2} \]

OpenStudy (anonymous):

In this case \(d=d_{pizza}\)

OpenStudy (anonymous):

While \(d_{target} = 16\)

OpenStudy (anonymous):

So is the final equation \[d \le3\times2^{-2}\]

OpenStudy (anonymous):

Nope...

OpenStudy (anonymous):

You have \[ |16-d|\leq 3\times 2^{-2} \]

OpenStudy (anonymous):

I mean shoot. One sec.

OpenStudy (anonymous):

\[d \le3\times1/4\]

OpenStudy (anonymous):

Is it that? (This is zero and negative exponents is the lesson)

OpenStudy (anonymous):

You can't ignoree the \(|16-d|\)

OpenStudy (anonymous):

But you got the exponent right.

OpenStudy (anonymous):

How come it is |16−d| What does it mean? Why is it in absolut value?

OpenStudy (anonymous):

It's in the absolute value because we want the difference to be positive.

OpenStudy (anonymous):

Because we need it to 16 or less. Right?

OpenStudy (anonymous):

For example, if \(d=18\) we don't want \(16-18=-2\), we want \(|16-18|=|-2|=2\)

OpenStudy (anonymous):

But it has to be within the targeted area of 16. Going over 16 wouldn't work.

OpenStudy (anonymous):

\[ | \dots |\leq a \quad\to\quad -a\leq \dots \leq a \]

OpenStudy (anonymous):

That doesn't seem like it would be an algerbra lesson. I'm in algebra right now.

OpenStudy (anonymous):

Well within could mean